# Outlier Gradient Analysis: Efficiently Identifying Detrimental Training Samples for Deep Learning Models

### v3

@@ C.5 Experiments on ResNet-18 Architecture
We also provide results for ResNet-18 [He et al., 2016] being used as the base model IN Table 9 instead of the ResNet-34 model. The overall performance of the ResNet-18 model is lower than ResNet-34 for all datasets and noise settings, since the ResNet-18 model has fewer residual connections than the ResNet-34 model. Moreover, it can be observed that outlier gradient analysis leads to improved performance post trimming, compared to the cross entropy baseline. Outlier gradient trimming is advantageous as a data selection strategy irrespective of the base model.
@@ C.6 Experiments on ImageNet
Although noisy label experiments have not been conducted on ImageNet [Deng et al., 2009], we decided to undertake a simple experiment on a subset of ImageNet. We created a subset of ImageNet containing 50000 images (50 images from each of the 1000 classes) as the training set, and flipped 40% of the corresponding image labels to create noisy labels (20 images from each class). The validation set is the same as ImageNet with 50000 images. We obtain results for performance on this set for a baseline ResNet-18 [He et al., 2016] model, DataInf, Gradient Tracing, iForest based outlier gradient analysis, as well as simple L1-norm and L2-norm thresholding based outlier gradient analysis. The models are trained for 10 epochs. In this limited experimental setting, we obtain the following results in Table 10 and find that outlier gradient analysis methods achieve competitive performance to other methods while being highly computationally efficient.
@@ C.7 Experiments on Other Noisy Learning Baselines
As we discussed previously, approaches for noisy learning can be categorized into (1) methods that either change the loss function or model architecture or (2) those that identify noisy samples and remove/relabel them for improving model performance [Algan and Ulusoy, 2021]. Since our approach belongs to the latter category, we only compared against other approaches from this category in the main paper. For completeness we now present results comparing our approach with some others in the former category for the ResNet-34 architecture and CIFAR-10N dataset. As can be seen in Table 11, outlier gradient analysis features in the top-2 performers compared to the other noisy learning baselines. We would like to emphasize that this is not an exhaustive list of baselines and noisy learning by adjusting the loss/model is not the primary focus of our work (but detecting detrimental samples is). Note that our algorithm could also be combined with approaches from this other category for additional gains.
@@
C.8 Experiments on Adapting Outlier Gradient Analysis to Validation/Test Set
We also conduct experiments for the distribution shift benchmark from the influence function work by Chhabra et al. [2024]. These experiments will showcase the applicability of outlier gradient analysis in adapting to a validation/test set distribution (instead of solely relying on the training set distribution). In Chhabra et al. [2024], three distribution shift scenarios are considered on the Folktables ACS-Income [Ding et al., 2021] dataset: time-shifted, location-shifted, and time+location-shifted. Essentially, in each of these settings, either the train/test distribution are time-shifted (e.g. 2014/2018), location-shifted (e.g. CA/MI), or both (e.g. 2014 & CA / 2018 & MI). We undertake the same experiments but using the OneClassSVM semi-supervised outlier analysis approach [Li et al., 2003] instead of iForest, L1/L2 norm, and provide the test set as inliers to correct the distribution of the training set influence estimation. Then, we utilize outlier gradient analysis for each setting, with results shown in Table 12. Our approach is highly adaptable to differing test/validation set distributions (concept drift) and can significantly outperform other baselines in this setting as well.
Table 12: Using OneClassSVM as the outlier analysis approach in the distribution shift experiments of Chhabra et al. [2024] on the Folktables ACS-Income dataset.
@@ C.8 Experiments on Adapting Outlier Gradient Analysis to Validation/Test Set
Appendix D Broader Impact and Limitations
Our work and proposed techniques aim to address the data-centric task of identifying detrimental samples. We improve upon the influence function analysis framework that is used to undertake this problem, but possesses deficiencies when applied to deep learning models. Enabling influence estimation for deep models allows practitioners to assess whether training samples are beneficial or detrimental to performance, and can make models more interpretable and performant. As we show through extensive experiments on multiple problem settings, our proposed outlier gradient analysis approach outperforms existing baselines and can augment model performance by identifying/trimming detrimental samples in a computationally efficient manner. As a result, our work paves the way for significant positive societal impact, especially with the increased adoption of larger and deeper neural networks such as LLMs. However, as with any work, there are limitations to our approaches that can be overcome in future work. For instance, it might be possible to derive specific outlier analysis algorithms that are computationally more efficient than iForest or norm thresholding, and significantly more performant. Another limitation that can be overcome is the further study and benchmarks for influence based analysis in LLMs– going beyond the datasets and approaches we used in this work. Finally, influence approaches can also be studied for generation tasks in vision based diffusion models.
Appendix E Code and Reproducibility
We provide our code, instructions, and implementation in an open-source repository: https://anonymous.4open.science/r/outlier-gradient-analysis/. The experiments were conducted on two separate Linux (Ubuntu 20.04.6 LTS) servers– the experiments of Sections 6 and 7 were conducted on NVIDIA GeForce RTX A6000 GPUs with 50GB VRAM running CUDA version 12.0 and all other experiments were conducted on an NVIDIA Tesla V100 with 32GB VRAM and CUDA version 11.4.

### v5

@@ C.5 Experiments on ResNet-18 Architecture
We also provide results for ResNet-18 (He et al., 2016) being used as the base model IN Table 9 instead of the ResNet-34 model. The overall performance of the ResNet-18 model is lower than ResNet-34 for all datasets and noise settings, since the ResNet-18 model has fewer residual connections than the ResNet-34 model. Moreover, it can be observed that outlier gradient analysis leads to improved performance post trimming, compared to the cross entropy baseline. Outlier gradient trimming is advantageous as a data selection strategy irrespective of the base model.
@@ C.6 Experiments on ImageNet
Although noisy label experiments have not been conducted on ImageNet (Deng et al., 2009), we decided to undertake a simple experiment on a subset of ImageNet. We created a subset of ImageNet containing 50000 images (50 images from each of the 1000 classes) as the training set, and flipped 40% of the corresponding image labels to create noisy labels (20 images from each class). The validation set is the same as ImageNet with 50000 images. We obtain results for performance on this set for a baseline ResNet-18 (He et al., 2016) model, DataInf, Gradient Tracing, iForest based outlier gradient analysis, as well as simple L1-norm and L2-norm thresholding based outlier gradient analysis. The models are trained for 10 epochs. In this limited experimental setting, we obtain the following results in Table 10 and find that outlier gradient analysis methods achieve competitive performance to other methods while being highly computationally efficient.
@@ C.7 Experiments on Other Noisy Learning Baselines
As we discussed previously, approaches for noisy learning can be categorized into (1) methods that either change the loss function or model architecture or (2) those that identify noisy samples and remove/relabel them for improving model performance (Algan & Ulusoy, 2021). Since our approach belongs to the latter category, we only compared against other approaches from this category in the main paper. For completeness we now present results comparing our approach with some others in the former category for the ResNet-34 architecture and CIFAR-10N dataset. As can be seen in Table 11, outlier gradient analysis features in the top-2 performers compared to the other noisy learning baselines. We would like to emphasize that this is not an exhaustive list of baselines and noisy learning by adjusting the loss/model is not the primary focus of our work (but detecting detrimental samples is). Note that our algorithm could also be combined with approaches from this other category for additional gains.
@@
C.8 Comparison with GEX (Kim et al., 2024) and TRAK (Park et al., 2023)
We also compare the performance of our outlier gradient analysis methods with GEX (Kim et al., 2024) and TRAK (Park et al., 2023) , two new influence function methods. As mentioned in the main paper, we were able to obtain results for CIFAR-10N but obtained out-of-memory (OOM) errors for CIFAR-100N. This computational memory overhead highlights the shortcomings of these approaches. Furthermore, the results on CIFAR-10N for all three noise settings are shown in Table 12. As can be seen, our outlier gradient analysis approaches outperform these new influence function baselines for the detrimental data identification task.
Table 12: Performance comparison of our outlier gradient analysis methods with GEX (Kim et al., 2024) and TRAK (Park et al., 2023) influence function baselines.
Method CIFAR-10N (Aggregate) CIFAR-10N (Random) CIFAR-10N (Worst)
GEX (Kim et al., 2024) 90.67 89.13 80.30
TRAK (Park et al., 2023) 91.73 90.07 83.52
Outlier Gradient (L1) 91.86 90.66 84.20
Outlier Gradient (L2) 92.21 90.25 82.99
Outlier Gradient (iForest) 91.36 90.20 83.72
C.9 Experiments on Adapting Outlier Gradient Analysis to Validation/Test Set
We also conduct experiments for the distribution shift benchmark from the influence function work by (Chhabra et al., 2024). These experiments will showcase the applicability of outlier gradient analysis in adapting to a validation/test set distribution (instead of solely relying on the training set distribution). In (Chhabra et al., 2024), three distribution shift scenarios are considered on the Folktables ACS-Income (Ding et al., 2021) dataset: time-shifted, location-shifted, and time+location-shifted. Essentially, in each of these settings, either the train/test distribution are time-shifted (e.g. 2014/2018), location-shifted (e.g. CA/MI), or both (e.g. 2014 & CA / 2018 & MI). We undertake the same experiments but using the OneClassSVM semi-supervised outlier analysis approach (Li et al., 2003) instead of iForest, L1/L2 norm, and provide the test set as inliers to correct the distribution of the training set influence estimation. Then, we utilize outlier gradient analysis for each setting, with results shown in Table 13. Our approach is highly adaptable to differing test/validation set distributions (concept drift) and can significantly outperform other baselines in this setting as well.
Table 13: Using OneClassSVM as the outlier analysis approach in the distribution shift experiments of (Chhabra et al., 2024) on the Folktables ACS-Income dataset.
@@ C.9 Experiments on Adapting Outlier Gradient Analysis to Validation/Test Set
Appendix D Code and Reproducibility
We provide our code, instructions, and implementation in an open-source repository: https://github.com/anshuman23/outlier-gradient-analysis. The experiments were conducted on two separate Linux (Ubuntu 20.04.6 LTS) servers– the experiments of Sections 6 and 7 were conducted on NVIDIA GeForce RTX A6000 GPUs with 50GB VRAM running CUDA version 12.0 and all other experiments were conducted on an NVIDIA Tesla V100 with 32GB VRAM and CUDA version 11.4.